# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v3

@@ Abstract
We derive formulae that resum, at a given order in the limit of small frequency, the infinite series of Post-Minkowkian (small gravitational coupling) or Post-Newtonian (small velocities) corrections to the gravitational waveform produced by particles moving along a general (open or closed) trajectory in Schwarzschild geometry. Specifying to the case of circular orbits, we compute the wave form and the energy flux to order 30PN, and compare it against the available results in the literature. The results are based on a novel hypergeometric representation of the solutions of the Heun equation (and its confluence), that leads to a mathematical proof of the Heun connection formula.
@@ 1 Introduction
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion Thorne:1980ru ; Damour:1982wm ; Blanchet:2013haa and perturbation theory Teukolsky:1973ha ; Poisson:1993vp ; Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk . The latter, which applies to EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
@@
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms, leading to an infinite tower of tails and tails of tails non-local interactions Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bianchi:2021mft ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a SU(2)2 quiver gauge theory. Following this path the infinite tower of logarithmic terms have been resummed into exponentials.
More recently, multipole Bini:2023fiz ; Brandhuber:2023hhy ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Herderschee:2023fxh ; Georgoudis:2023eke ; Georgoudis:2023lgf ; Georgoudis:2023ozp ; Elkhidir:2023dco ; Bini:2024ijq ; Bini:2024rsy ; Georgoudis:2024pdz ; Alessio:2024onn ; Brunello:2024ibk and perturbation theory Fucito:2024wlg methods have been applied to the study of scattering processes (open trajectories) in gravity Kovacs:1977uw ; Kovacs:1978eu , a field where we can largely benefit from the impressive developments and refinements of scattering amplitude techniques in the last few years (see DiVecchia:2023frv for a review). These systems provide a richer set up, where the couplings of the PM and PN expansions are no longer linked to each other. For example, in the scattering amplitude language, the PM expansion corresponds to the expansion in loops. At a given order in the coupling, the amplitude is exact in the velocity, i.e. it resums the infinite PN series. To reproduce this result in perturbation theory, one has to resum the double expansion in one of the two variables.
For simplicity we focus on the geodesic motion in a Schwarzschild geometry and parametrize the mass, wavefrequency and radial distance by the two dimensionless combinations $z=2M/r$ and $y=2i\omega r$, assuming $x=4{\rm i}\omega M$ is small. We will show that, at order $x^{k}$, the dependence of the waveform on the variable $y$ or $z$ can be resummed and written in terms of a single hypergeometric function and its derivative, with coefficients given by two polynomials of degree $k$, that can be determined recursively. The two limits where $x$ is taking to be small keeping finite $y$ or $z$ lead to two alternative hypergeometric representations of the confluent Heun function, describing the PM and PN expansions respectively.
We observe that, the hypergeometric representation of the confluent Heun function we derived here is similar in spirit to the well studied MST solution Mano:1996vt . The crucial difference between the two, is that while the MST ansatz involves an infinite sum of hypergeometric functions, our solution is written in terms of a single hypergeometric function and its first derivative.
@@
This leads to a mathematical proof of the Heun connection formulae originally found in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten . The whole non-triviality of the Heun equation (and its confluences) is codified in a single function, $a(\ell)$, the Seiberg-Witten (SW) period of the underlying ${\cal N}=2$ supersymmetric gauge theory (see also Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe ). This function is related to the so called "renormalized angular momentum" in the MST context Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk or "anomalous dimension" in the effective field theory context Goldberger:2009qd ; Trestini:2023wwg ; Almeida:2021jyt ; Edison:2023qvg . The Heun connection formula has been recently applied in a variety of physical contexts, including the study of QNMs for gravity backgrounds (BHs, fuzzballs, D-branes), correlators in thermal field theories, holography, cosmological perturbations and scattering amplitudes in gravity Bianchi:2021xpr ; Bianchi:2021mft ; Fioravanti:2021dce ; Bianchi:2022wku ; Dodelson:2022eiz ; Dodelson:2022yvn ; Imaizumi:2022qbi ; Fioravanti:2022bqf ; Imaizumi:2022dgj ; Bianchi:2022qph ; Dodelson:2023vrw ; Bianchi:2023rlt ; Bianchi:2023sfs ; Giusto:2023awo ; Aminov:2023jve ; BarraganAmado:2023apy ; Lei:2023mqx ; BarraganAmado:2023wxt ; Fucito:2023afe ; Bautista:2023sdf ; Ge:2024jdx ; Cipriani:2024ygw ; Arnaudo:2024rhv ; Bianchi:2024vmi ; Bianchi:2024mlq ; DiRusso:2024hmd ; Bianchi:2024rod ; Aminov:2024aan ; BarraganAmado:2024tfu ; Bautista:2024emt ; Liu:2024eut ; Silva:2025khf ; Aprile:2025hlt ; Bhatta:2025kil .
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. Finally, we remark, that even if the main motivation for resuming PM and PN series relies on the study of open orbits, our hypergeometric representations provide an algorithm which is very efficient when implemented in a Mathematica code to describe PN expansions of bounded orbits. Indeed, here we compute the gravitational waveform and the energy flux, in the case of circular orbits, up to the 30th PN order [^2] The ancillary mathematica files coming with this paper should help the reader to check our computations.. This result will be tested against the 22nd PN order computation available in literature Fujita:2012cm , and results based on MST and numerical methods from the Black Hole Perturbation Toolkit (BHTK) BHPToolkit . The analytic expressions of all ingredients contributing to the 30PN expansion of the wave form and energy flux are collected in ancillary mathematical files. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. In Section 5 we extend the results to include the case of the non-confluent Heun function. Section 6 contains our conclusions.
@@ 2 The confluent Heun equation
$g^{\prime\prime}(t)+\left(\frac{\beta_{0}}{t}+\frac{\beta_{1}}{t-1}+\lambda\right)g^{\prime}(t)+\frac{\alpha_{1}t-\alpha_{0}}{t(t-1)}g(t)=0$ (1)
@@ 2.1 Confluent Heun-Gauge dictionary
The confluent Heun equation can be viewed as the quantum version of the Seiberg-Witten (SW) curve governing the dynamics of a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with three massive flavours ($m_{i}$)transforming in the fundamental representation of the gauge group (with gauge coupling $x$) Bianchi:2021mft ; Bianchi:2021xpr living on a curved Nekrasov-Shatashvilli background Nekrasov:2009rc . This suggests a a change of variables in the Heun equation (1) in terms of the parameters of the SW quantum curve and its Coulomb branch parameter $u$
$\alpha_{0}$ $=$ $u-{\textstyle\frac{1}{4}}+m_{1}-m_{1}^{2}+(m_{1}+m_{3}-1)x\qquad,\qquad\alpha_{1}=\left(m_{1}+m_{3}-1\right)x$
$\beta_{0}$ $=$ $1-m_{1}+m_{2}\qquad,\qquad\beta_{1}=1-m_{1}-m_{2}\qquad,\qquad\lambda=-x$ (3)
which will simplify the computations which we are going to perform later in the paper. Defining
$G_{1}(z)=z^{-{\alpha_{1}\over\lambda}}g({\textstyle\frac{1}{z}})$ (4)
the Heun equation (1) can be put into the form [^3] Alternatively, equation (8) can be written in the normal Schrödinger like form $\Psi^{\prime\prime}(z)+Q_{\rm gauge}(z)\Psi=0\qquad,\qquad G_{1}(z)=e^{{\frac{x}{2z}}}z^{m_{3}-1}(1-z)^{m_{1}+m_{2}-1\over 2}\Psi(z)$ (5) with $Q_{\rm gauge}={-}\frac{x^{2}}{4z^{4}}{+}\frac{xm_{3}}{z^{3}}{+}\frac{1{-}\left(m_{1}{-}m_{2}\right)^{2}}{4(z{-}1)z}+\frac{1{-}\left(m_{1}{+}m_{2}\right){}^{2}}{4(z-1)^{2}z}+\frac{u{-}{\textstyle\frac{1}{4}}{+}{\textstyle\frac{x}{2}}\left(m_{1}{+}m_{2}{-}1\right)}{(z{-}1)z^{2}}$ (6) Notice that also the conformal Ward identity of Liouville theory, the so called BPZ equation, is a confluent Heun equation. The normal form of such equation matches again with (5) and (6)after the identifications $m_{1}=p_{0}-k_{0}\qquad,\qquad m_{2}=-p_{0}-k_{0}\qquad,\qquad m_{3}=c\qquad,\qquad x=-x_{\rm CFT}$ (7) with $k_{0}$, $p_{0}$, $c$ and $x_{\rm CFT}$ the CFT variables in Consoli:2022eey ; Fucito:2023afe .
$G_{1}^{\prime\prime}(z)+G_{1}^{\prime}(z)\left(\frac{1{-}m_{1}{-}m_{2}}{z{-}1}{-}\frac{2\left(m_{3}{-}1\right)}{z}{+}\frac{x}{z^{2}}\right)$ (8)
$\qquad{+}G_{1}(z)\left(\frac{{\textstyle\frac{1}{4}}{-}u{+}m_{3}(m_{3}{-}1)}{z^{2}}{+}\frac{u+\frac{3}{4}-m_{1}-m_{2}-m_{3}+m_{1}m_{2}+m_{1}m_{3}+m_{2}m_{3}}{(z{-}1)z}\right)=0$
where now the irregular point is located at the origin. Alternatively, one writes
$G_{0}(z)=G_{1}({\textstyle\frac{x}{z}})$ (9)
and setting $y=x/z$ the confluent Heun equation (8) becomes
$G_{0}^{\prime\prime}(y){+}G_{0}^{\prime}(y)\left(\frac{1{-}m_{1}{-}m_{2}}{y{-}x}{+}\frac{m_{1}{+}m_{2}{+}2m_{3}{-}1}{y}{-}1\right)$ (10)
$\qquad\qquad{+}G_{0}(y)\left(\frac{u{+}{\textstyle\frac{3}{4}}{+}\delta}{(x-y)y}{+}\frac{1{+}\delta{+}m_{3}(m_{3}{-}1)}{y^{2}}\right)=0$
@@
where $\delta=m_{1}m_{2}+m_{1}m_{3}+m_{2}m_{3}-m_{1}-m_{2}-m_{3}$. We will use both descriptions of the confluent Heun equation, and we will refer to them as CHE. In particular, we will consider the expansions of the two equations in the limit where $x$ is taken small but keeping finite either $z$ or $y=x/z$. The two expansions will correspond in the gravity context to the Post-Minkowskian and Post-Newtonian approximations, so we will refer to $G_{0}$ and $G_{1}$ as the PM and PN representations of the confluent Heun function respectively. In particular, setting $x\to 0$, (10) and (8) reduce to hypergeometric equations with solutions
@@
with $\alpha=\pm$ labeling the two independent solutions,
@@
and $a\underset{x\to 0}{\approx}\sqrt{u}$. On the other hand, at order $k$ in $x$, we will look for solutions of the form
@@
one can determine the coefficients $c^{PM}_{ij}(\alpha a)$, $c^{PN}_{ij}(\alpha a)$ and $u_{i}(a)$ by solving the CHE equations order by order in $x$. We notice that $a$ codify the monodromy of the solutions around the irregular point $z=0$. Moreover, since $G^{1}_{\alpha}(z)$ and $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ are solutions of the same equation and have the same monodromy around $z=0$ they should be proportional to each other. We conclude therefore that in the overlap region, the so called near zone, where $x\ll z\ll 1$
@@
The main result of this Section is that, the expansions PMk and PNk in (2.1) can be exactly resumed and written in terms of a single hypergeometric function. This will provide us with a first principle proof of the connection formulae for the confluent Heun functions, where the whole non-triviality of the Heun equation is codified into a single function $a(u)$.
@@
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations made out of the four coefficients $B_{\pm}$, $D_{\pm}$. We will derive connection formulae relating the coefficients $B_{\pm}$ to $D_{\pm}$ and viceversa.

### v4

@@ Abstract
We derive formulae that resum, at a given order in the soft limit, the infinite series of Post-Minkowskian (small gravitational coupling) or Post-Newtonian (small velocities) corrections to the gravitational waveform produced by particles moving along a general (open or closed) trajectory in the Schwarzschild geometry in the probe limit. Specifying to the case of circular orbits, we compute the waveform and the energy flux to order 30PN, and compare it against the available results in the literature. Our results are based on a novel hypergeometric representation of the solutions of the Heun equation (and its confluence), that leads to a simple mathematical proof of the Heun connection formula.
@@ 1 Introduction
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion Thorne:1980ru ; Damour:1982wm ; Blanchet:2013haa and perturbation theory Teukolsky:1973ha ; Poisson:1993vp ; Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk . The latter, valid in the probe limit of the binary, the so called EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
@@
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms, leading to an infinite tower of tails and tails of tails non-local interactions Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bianchi:2021mft ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms, in the probe limit, were proposed in terms of the instanton partition function of a SU(2)2 quiver gauge theory. Following this path the infinite tower of logarithmic terms have been resummed into exponentials.
More recently, multipole methods Bini:2023fiz ; Bini:2024ijq ; Bini:2024rsy , scattering amplitude techniques Jakobsen:2021smu ; Mougiakakos:2021ckm ; Cristofoli:2021vyo ; DiVecchia:2023frv ; Brandhuber:2023hhy ; Herderschee:2023fxh ; Elkhidir:2023dco ; Georgoudis:2023lgf ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Georgoudis:2023eke ; Georgoudis:2023ozp ; Georgoudis:2024pdz ; Brunello:2024ibk ; Alessio:2024onn and black hole perturbation theory Fucito:2024wlg have been applied to the study of gravitational scattering processes, fostering a fruitful interplay and mutual cross-checking among these complementary approaches. Within the quantum field theory framework, the PM series describes the expansion in loops (weak coupling) of the scattering amplitude keeping finite the relative velocity of the binary. Using perturbation theory, we will show that at any fixed order in the gravitational coupling, the entire series of PN corrections in the probe limit of the gravitational waveform can be resummed and expressed in terms of a single hypergeometric function. Analogous results hold for the PN expansion itself, where the exact dependence on the gravitational coupling corrections is determined at any given order in the small-velocity limit.
@@
The PM and PN series provide novel hypergeometric representations of the confluent Heun function along the lines of the well studied MST solutions Mano:1996vt . The main difference between the two, is that the MST ansätze involves an infinite sum of hypergeometric functions, while the solution we put forward here is written in terms of a single hypergeometric function and its first derivative. The crucial advantage of this representation is that it provides a simple mathematical proof of the Heun connection formulae originally found in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten . The connection matrix coefficients are given as ratios of gamma functions with arguments depending on a single (but highly non-trivial) function codifying the holonomies of the confluent Heun function around the singular point. This function, known as the "renormalized angular momentum" in the MST context Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk , can be related to the so called Seiberg-Witten period of a ${\cal N}=2$ supersymmetric gauge theory Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe or the "anomalous dimension" in the effective field theory description of the gravitational wave Goldberger:2009qd ; Trestini:2023wwg ; Almeida:2021jyt ; Edison:2023qvg . The Heun connection formula has been recently applied in a variety of physical contexts, including the study of QNMs for gravity backgrounds (BHs, fuzzballs, D-branes), correlators in thermal field theories, holography, cosmological perturbations and scattering amplitudes in gravity Bianchi:2021xpr ; Bianchi:2021mft ; Fioravanti:2021dce ; Bianchi:2022wku ; Dodelson:2022eiz ; Dodelson:2022yvn ; Imaizumi:2022qbi ; Fioravanti:2022bqf ; Imaizumi:2022dgj ; Bianchi:2022qph ; Dodelson:2023vrw ; Bianchi:2023rlt ; Bianchi:2023sfs ; Giusto:2023awo ; Aminov:2023jve ; BarraganAmado:2023apy ; Lei:2023mqx ; BarraganAmado:2023wxt ; Fucito:2023afe ; Bautista:2023sdf ; Ge:2024jdx ; Cipriani:2024ygw ; Arnaudo:2024rhv ; Bianchi:2024vmi ; Bianchi:2024mlq ; DiRusso:2024hmd ; Bianchi:2024rod ; Aminov:2024aan ; BarraganAmado:2024tfu ; Bautista:2024emt ; Liu:2024eut ; Silva:2025khf ; Aprile:2025hlt ; Bhatta:2025kil .
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions (and its confluences) from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. Finally, we remark, that even if the main motivation for resumming PM and PN series relies on the study of open orbits, our hypergeometric representations provide an algorithm which is very efficient when implemented in a Mathematica code to describe the PN expansions of bounded orbits. Indeed, here we compute the gravitational waveform and the energy flux, in the case of circular orbits, up to 30 PN orders. This result will be tested against the computation at 22 PN orders available in literature Fujita:2012cm and the numerical methods from the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST approach. The analytic expressions of all ingredients contributing to the 30PN expansion of the wave form and energy flux are collected in ancillary mathematical files. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. Section 5 contains our conclusions. In Appendix A we derive the PM and PN expansions of the non-confluent Heun function.
Conventions
The second order ordinary differential equation called the confluent Heun equation is introduced in (6). It has two regular singularities at $z=1,\infty$ and an irregular one at $z=0$. The equation is specified by five complex parameters dubbed here as $\{m_{1},m_{2},m_{3},u,x\}$. The expansion of the general solution of this equation in the ”soft limit” $x\ll 1$ is derived in Section 2. These results apply to any physical problem governed by an equation of confluent Heun type. The focus of this paper is on the study of gravitational wave forms in the Schwarzschild geometry. The Heun-gravity dictionary in this case reads
$x$ $=$ $4{\rm i}\omega M\qquad,\qquad y=2{\rm i}\omega r\qquad,\qquad z={2M\over r}$
$m_{1}$ $=$ ${\textstyle\frac{x}{2}}\quad,\quad m_{2}=m_{3}^{*}=-2+{\textstyle\frac{x}{2}}\quad,\quad u=(\ell+{\textstyle\frac{1}{2}})^{2}+{\textstyle\frac{5x}{2}}+{\textstyle\frac{x^{2}}{4}}$ (1)
with $y=x/z$, $M$ the mass of the BH, $\omega$ and $\ell$ the frequency and orbital number of the gravitational wave and $r$ the radial distance. Along the paper we set $G=1$ and use the mostly plus signature ($+$)(-$+$)( - + + + ).
@@ 2 The confluent Heun equation
${\rm HeunC}^{\prime\prime}(t)+\left(\frac{\beta_{0}}{t}+\frac{\beta_{1}}{t-1}+\lambda\right){\rm HeunC}^{\prime}(t)+\frac{\alpha_{1}t-\alpha_{0}}{t(t-1)}{\rm HeunC}(t)=0$ (2)
@@ 2.1 Confluent Heun-Gauge dictionary
The confluent Heun equation can be viewed as the quantum version of the Seiberg-Witten (SW) curve governing the dynamics of a ${\cal N}=2$ supersymmetric SU(2) gauge theory with three flavours of masses $m_{i}$, transforming in the fundamental representation of the gauge group Bianchi:2021mft ; Bianchi:2021xpr living on a curved Nekrasov-Shatashvili background Nekrasov:2009rc . The quantum curve can be written in the form [^2] Equation (6) can be written in the normal Schrödinger like form $\Psi^{\prime\prime}(z)+Q_{\rm gauge}(z)\Psi=0\qquad,\qquad G_{1}(z)=e^{{\frac{x}{2z}}}z^{m_{3}-1}(1-z)^{m_{1}+m_{2}-1\over 2}\Psi(z)$ (4) with $Q_{\rm gauge}={-}\frac{x^{2}}{4z^{4}}{+}\frac{xm_{3}}{z^{3}}{+}\frac{1{-}\left(m_{1}{-}m_{2}\right)^{2}}{4(z{-}1)z}+\frac{1{-}\left(m_{1}{+}m_{2}\right){}^{2}}{4(z-1)^{2}z}+\frac{u{-}{\textstyle\frac{1}{4}}{+}{\textstyle\frac{x}{2}}\left(m_{1}{+}m_{2}{-}1\right)}{(z{-}1)z^{2}}$ (5)
$G_{1}^{\prime\prime}(z){+}G_{1}^{\prime}(z)\left(\frac{1{-}m_{1}{-}m_{2}}{z{-}1}{-}\frac{2\left(m_{3}{-}1\right)}{z}{+}\frac{x}{z^{2}}\right){+}G_{1}(z)\left(\frac{{\textstyle\frac{1}{4}}{-}u{+}m_{3}(m_{3}{-}1)}{z^{2}}{+}\frac{u{+}\frac{3}{4}{+}\delta}{(z{-}1)z}\right)=0$ (6)
with $\delta=m_{1}m_{2}{+}m_{1}m_{3}{+}m_{2}m_{3}{-}m_{1}{-}m_{2}{-}m_{3}$, $x$ the gauge coupling and $u$ the Coulomb branch parameter. (2) and (6) are related by the Heun-gauge dictionary
$G_{1}(z)$ $=$ $z^{-{\alpha_{1}\over\lambda}}{\rm HeunC}({\textstyle\frac{1}{z}})\qquad,\qquad t=z^{-1}\qquad,\qquad\lambda=-x$
$\alpha_{0}$ $=$ $u-{\textstyle\frac{1}{4}}+m_{1}(1-m_{1})+(m_{1}+m_{3}-1)x\qquad,\qquad\alpha_{1}=\left(m_{1}+m_{3}-1\right)x$
$\beta_{0}$ $=$ $1-m_{1}+m_{2}\qquad,\qquad\beta_{1}=1-m_{1}-m_{2}$ (7)
Equivalently, the confluent Heun equation can be written as a differential equation in the variable $y=x/z$, taking the form
$G_{0}^{\prime\prime}(y){+}G_{0}^{\prime}(y)\left(\frac{1{-}m_{1}{-}m_{2}}{y{-}x}{+}\frac{m_{1}{+}m_{2}{+}2m_{3}{-}1}{y}{-}1\right){+}G_{0}(y)\left(\frac{u{+}{\textstyle\frac{3}{4}}{+}\delta}{(x-y)y}{+}\frac{1{+}\delta{+}m_{3}(m_{3}{-}1)}{y^{2}}\right)=0$ (8)
with
$g(x)G_{0}(y)=G_{1}({\textstyle\frac{x}{y}})=({\textstyle\frac{y}{x}})^{{\alpha_{1}\over\lambda}}{\rm HeunC}({\textstyle\frac{y}{x}})$ (9)
@@
and $g(x)$ an arbitrary constant. We will use both descriptions (6) and (8) of the confluent Heun equation, and we will refer to them as CHE. In particular, we will focus on the soft (or large wavelength) limit where $x$ is taken small keeping finite either $z$ or $y$. In the gravity context, the two limits correspond to the PM and PN expansions of the confluent Heun solution, so we will refer to $G_{0}$ and $G_{1}$ as the PM and PN representations of the confluent Heun function respectively. In particular, sending $x\to 0$, (8) and (6) reduce to hypergeometric equations with solutions
@@
with $\alpha=\pm$ labelling the two independent solutions,
@@
and $a\underset{x\to 0}{\approx}\sqrt{u}$. On the other hand, at order $k$ in $x$,  [^3] This is the only expansion that will appear in our formulae. With an abuse of notation we will define our results a PN expansion at order $k$ even if other contributions to our final formulae will be given in closed form. we will look for solutions of the form
@@
one can determine the coefficients $c^{PM}_{ij}(a)$, $c^{PN}_{ij}(a)$ and $u_{i}(a)$ by solving the CHE equations order by order in $x$. We notice that $a$ codifies the monodromy of the solutions around the irregular point $z=0$. Moreover, since $G^{1}_{\alpha}(z)$ and $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ are solutions of the same equation and have the same monodromy around $z=0$ they should be proportional to each other. We conclude therefore that in the overlap region, the so called near zone, where $x\ll z\ll 1$
@@
The main result of this Section is that, the series PMk and PNk in (2.1) can be exactly resummed and written in terms of a single hypergeometric function. This will provide us with a first principle proof of the connection formulae for the confluent Heun functions, where the whole non-triviality of the Heun equation is codified into a single function $a(u)$.
@@
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations out of the four coefficients $B_{\pm}$, $D_{\pm}$. We will derive connection formulae relating the coefficients $B_{\pm}$ to $D_{\pm}$ and viceversa.